OpenStudy (anonymous):

What is the 12th term of the arithmetic sequence where a1 = 5 and a6 = 20 ?

OpenStudy (anonymous):

38?

OpenStudy (anonymous):

thats what i got

OpenStudy (anonymous):

in order to find the common difference we use formula l=a1+(n-1)d here a1=5, l=20 n=6 putting these all in formula we get d=3 and now solving for the 12th term a12=5+(12-1)3 it implies that 12th term is 38

OpenStudy (anonymous):

5 + (12-1)((20-5)/5) = 38